OpenStudy (vshiroky):

Find the missing coordinate to complete the ordered pair solution to the given linear equation. 2y+5x=-10 A. (-2,__) B. (__,5)

OpenStudy (vshiroky):

@jim_thompson5910 Can you please help me I'm sure I will have more after this I just want to make sure I know what I am doing.

jimthompson5910 (jim_thompson5910):

A) The ordered pair (-2, ___) means that we know for sure that x = -2 So plug in x = -2 and solve for y 2y+5x=-10 2y+5(-2)=-10 2y-10=-10 2y-10+10=-10+10 ... Add 10 to both sides. 2y+0= 0 2y = 0 y = 0/2 y = 0 So if x = -2, then y = 0 Therefore, (-2,___) turns into (-2,0)

OpenStudy (vshiroky):

I believe its A. (-2,0) and B. (-4,5)

jimthompson5910 (jim_thompson5910):

B) 2y+5x=-10 2(5)+5x=-10 10+5x = -10 5x = -10-10 5x = -20 x = -20/5 x = -4 So the other ordered pair is (-4,5) You are correct on both parts. Nice job.

OpenStudy (vshiroky):

Great!! Thank you. I'm sure I will have more haha. I just like to verify things

jimthompson5910 (jim_thompson5910):

It looks like you definitely know what you're doing, which is great. Keep up the good work.